OpenStudy (anonymous):

What is the exact area of a circle with a diameter of 18 feet? Just want the answer

OpenStudy (anonymous):

you look like bieber

OpenStudy (anonymous):

lol ok

OpenStudy (anonymous):

divide 18 by 2.

OpenStudy (anonymous):

then use your answer at plug it into r x r x 3.14

OpenStudy (campbell_st):

for the exact value... just write it in terms of pi... \[A = 9^2 \times \pi\]

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

254.469004941 or 81(pi)

OpenStudy (campbell_st):

a decimal answer is not exact....

OpenStudy (yanasidlinskiy):

That's what i was just going to do @campbell_st